Using Trioleoyl glycerol (trioleoin) determine and present the following information regarding the animal cell’s ability to obtain energy from the triglyceride molecule.
a. Total number of ATP’s that could theoretically be obtained from a complete breakdown (hydrolysis of ester bonds and beta oxidation of each fatty acid [FA]) of the triglyceride. Be sure to follow the molecules through all pertinent aspects of glycolysis, Kreb’s cycle, and the electron transport chain (ETC). Also, assume that the cell will obtain 3 ATP’s for each NADH entering the ETC and will obtain 2 ATP’s for each FADH2 entering the ETC.
b. The structural formula for each FA and the total number of carbons in each FA obtained from the hydrolysis of the ester bonds.
c. Total number of acetyl CoA’s produced through beta oxidation.
d. Total number of NADH’s produced from the oxidation of the triglyceride (be sure to consider NADH gained or lost from beta oxidation)
e. Total number of FADH2’s produced from the oxidation of the triglyceride (be sure to consider the FADH2’s gained or lost from beta oxidation)
f. The percent increase or decrease over the theoretical 36 ATP’s from 1 molecule of glucose.
Answer #1

B.

CH2-O-co-CH3(CH2)7CH=CH(CH2)7COOH
|
CHO-co-CH3(CH2)7CH=CH(CH2)7COOH
|
CH2-O-co-CH3(CH2)7CH=CH(CH2)7COOH

Oleic acid has one double bond and it is 18 carbon chain. In Beta oxidation one less FADH2 is produced.
In these 9 aceyl coA, 7 FADH2 and 8 NADH are produced by beta oxidation.

Activation of the fatty acid to Acyl CoA = -2 ATP
7 FADH2 = 7 X 2 ATP = 14 ATP
8 NADH = 8 X 3 ATP = 24 ATP

In krebs cycle
Acetyl CoA oxidation in the Krebs Cycle = 3NADH, 1FADH2, 1GTP
1 FADH2 = 1 X 2 ATP = 2 ATP
3 NADH = 3 X 3 ATP = 9 ATP
1GTP                       = 1ATP
1 Acetyl COA in TCA cycle provides 12 ATP
9 Acetyl CoA x 12 ATP/Acetyl CoA = 108 ATP
Total oleic acid ATP produces = 14+24+108-2 = 144 ATP

So for 3 oleic acid gives = 144 X 3 = 432 ATP.

1 Glycerol gives = 11 + 5 + 19 = 35 ATP

Total = 432 + 35 = 467 ATP produced from Trioleoyl glycerol.


F. 36/467 X 100 = 1300%